An all equity firm has a market value of $400,000 with 40,000 shares of stock. They...

  1. An all equity firm has a market value of $400,000 with 40,000 shares of stock. They are looking at changing their structure to include $100,000 in debt.

1. What is the current share price?

2. How many shares will $100,000 buy?

3. How many shares will remain after the capital change?

1. 10

2. 10,000

3. 40,000

1. 40

2. 2,500

3. 37,500

1. 10

2. 10,000

3. 30,000

1. 10

2. 10,000

3. 0
Answer #1

1. Current Share Price =Market Value/Number of shares =400000/40000 =10
2. Number of shares =Debt/Share Price =100000/10 =10000
3. Number of Shares after buy back =Number of shares initially -Number of shares used for debt =40000-10000 =30000

Option c is correct option